OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (michele_laino):

the length of your vector is: \[\Large \left\| u \right\| = \sqrt {{{\left( { - 4} \right)}^2} + {{\left( { - 3} \right)}^2}} = \sqrt {16 + 9} = 5\] now you have to divide your vector by its length

OpenStudy (michele_laino):

namely, what is: \[\Large \frac{u}{{\left\| u \right\|}}\]

OpenStudy (anonymous):

how do i put it in component form?? @Michele_Laino

OpenStudy (michele_laino):

it is simple: \[\Large \frac{u}{{\left\| u \right\|}} = \frac{{ < - 4, - 3 > }}{5} = < \frac{{ - 4}}{5},\frac{{ - 3}}{5} > \]

OpenStudy (michele_laino):

that procedure above is the division of a vector by a real number, which is always possible when, as in our case, that real number is different from zero

OpenStudy (anonymous):

ohhh okay thank you @Michele_Laino